OpenStudy (anonymous):

Let f(x) = x - 2 and g(x) = x2 - 7x - 9. Find f(g(-1)). -21 -3 3 21

OpenStudy (anonymous):

The term f(g(-1)) means that you plug in -1 in place of x in g(x) and then you take g(-1) and plug that in place of x for f(x). g(-1) = (-1)^2 - 7(-1) - 9 g(-1) = -2+7-9 g(-1) = -1 now you plug -1 (which is g(-1)) into f(x) in place of x. f(-1) = -1 - 2 f(-1) = -3

OpenStudy (anonymous):

You made it seem the numbers at the end were the answers...

OpenStudy (anonymous):

wait im so sorry you are right im wrong

OpenStudy (anonymous):

i was looking at the wrong thing, thankyou so much. @JjJustin